OpenStudy (billyjean):

ABC has side lengths 20,24, and 31. Do the side lengths form a Pythagorean triple?

OpenStudy (anonymous):

Does A^2 + B^2 = C^2?

OpenStudy (billyjean):

yes

OpenStudy (here_to_help15):

\( 20^2 * 24^2= 31^2 ? \)

OpenStudy (mathmath333):

here \(C>B,A\)

OpenStudy (anonymous):

Then t forms a triple

OpenStudy (here_to_help15):

So 20 * 20 =400 and 24*24= 576

OpenStudy (anonymous):

But A^2 + B^2 does not equal C^2

OpenStudy (mathmath333):

it should be this \(20^2 \color{red}{+} 24^2= 31^2 ?\)

OpenStudy (here_to_help15):

\(576+400= 976\)

OpenStudy (billyjean):

the answer choices are yes or no

OpenStudy (here_to_help15):

Oops didnt see that thanks @mathmath333

OpenStudy (here_to_help15):

Ok so \(A^2+B^2 = 976 \) does C^2 = 976 ? If not then it is no.

OpenStudy (here_to_help15):

So does \(31*31=976\) ?

OpenStudy (here_to_help15):

@billyjean ?

OpenStudy (billyjean):

yes it equals to 976

OpenStudy (here_to_help15):

No no no it does not lets multiply it again and see what it turns out to be.

OpenStudy (here_to_help15):